determine the point of intersection of each pair of functions
f(x)= -2x^-5x+20, g(x)=6x-1

Answer by srinivas(5)

to solve for point of intersection equate f(x) and g(x)
-2x^2-5x+20=6x-1
-2x^2-11x+21=0
2x^2+11x-21=0
2x^2+14x-3x-21=0
2x(+7)-3(x+7)=0
(2x-3)=0 (x+7)=0
therefore x=(3/2),-7